# Depth Pro: Sharp Monocular Metric Depth in Less Than a Second

### v1

@@ B.2 High-resolution alternatives
B.3 Training objectives
@@
Stage 1 training objectives. We first evaluate loss combinations for the first stage (Tab. 9).
HRWSI | Hypersim | Apolloscape | ||||||
Cond. | AbsRel | Log | AbsRel | Log | AbsRel | |||
1A | 0.166 | 82.1 | 0.083 | 0.259 | 75.4 | 0.156 | 0.339 | 45.6 |
1D | 0.138 | 85.1 | 0.077 | 0.246 | 78.4 | 0.128 | 0.424 | 60.6 |
1B | 0.156 | 83.3 | 0.078 | 0.249 | 77.3 | 0.152 | 0.300 | 47.3 |
1C | 0.150 | 83.7 | 0.074 | 0.235 | 79.9 | 0.084 | 0.235 | 75.6 |
Condition 1A only applies a mean absolute error loss to all datasets. For non-metric datasets, we use the scale-and-shift-invariant version. Condition 1B adds gradient losses to all synthetic datasets. We again use the scale-and-shift-invariant version for non-metric datasets. Following our observations from Sec. 3, we propose to apply an appropriate mean absolute error loss as in other conditions depending on a dataset being metric, but apply a scale-and-shift-invariant gradient loss irrespective of a dataset being metric or not (C). Although the intuition of applying a gradient loss is to sharpen the boundaries and improve edge-related metrics, it also improves other, non-boundary-related metrics. The key finding is that the gradient loss must be scale-and-shift-invariant, irrespective of the dataset being metric or not. Applying a regular gradient loss on metric datasets that early in training actually harms convergence.
Stage 2 training objectives.
HRWSI | Hypersim | Apolloscape | ||||||||||
@@
AbsRel | Log | AbsRel | Log | AbsRel | |||
2A | ✓ | ✓ | ✓ | ✓ | 0.149 | 83.6 | 0.072 | 0.235 | 81.3 | 0.092 | 0.303 | 72.9 |
2B | ✓ | ✓ | ✓ | 0.148 | 83.7 | 0.072 | 0.230 | 81.0 | 0.092 | 0.299 | 73.1 | |
2C | ✓ | ✓ | 0.150 | 83.7 | 0.072 | 0.235 | 80.8 | 0.091 | 0.300 | 73.2 | ||
2D | ✓ | 0.150 | 83.4 | 0.074 | 0.239 | 79.8 | 0.096 | 0.349 | 72.8 | |||
2E | 0.159 | 82.7 | 0.074 | 0.242 | 80.6 | 0.096 | 0.346 | 73.3 |
The second stage of our training curriculum focuses on sharpening depth boundaries. To that end, we only employ synthetic datasets due to their high quality ground truth. The obvious strategy for sharpening predictions is the application of gradient losses. We evaluate our combination of multiscale derivative-based losses in an ablation study. Condition 2A uses all of the losses, namely , ,, , and . Tab. 10. 2B removes the second-order loss . 2C further removes the squared first order losses . 2D removes all derivative-based losses. 2E applies the to all datasets. Removing improves results on Apolloscape. Our combination of 0th- to 2nd-order derivative losses (2A) performs best across metrics and datasets in aggregate (e.g., in terms of the average rank across metrics).
B.4 Full curricula
HRWSI | Hypersim | Apolloscape | ||||||
Cond. | AbsRel | Log | AbsRel | Log | AbsRel | |||
3A (Ours) | 0.149 | 83.6 | 0.072 | 0.235 | 81.3 | 0.092 | 0.303 | 72.9 |
3B | 0.148 | 83.9 | 0.073 | 0.245 | 81.3 | 0.095 | 0.292 | 72.1 |
3C | 0.153 | 83.6 | 0.166 | 0.386 | 37.1 | 0.586 | 0.712 | 0.5 |
Finally, we assess the efficacy of our complete training curriculum in comparison to alternatives. Condition 3A represents our two-stage curriculum. Condition 3B trains in a single stage and applies all the second-stage gradient losses throughout the whole training. Condition 3C reverses our two stages and represents the established strategy of pretraining on synthetic data first and fine-tuning with real-world data.
B.5 Focal length estimation
@@
DPT only | 60.0 |
ViT only | 74.4 |
DPT & ViT in series | 63.6 |
DPT & ViT in parallel | 78.2 |
Controlled evaluation of network structures. We evaluate a number of choices for the focal length estimation head and report results in Tab. 12. The models are evaluated on 500 images randomly sampled from Flickr (Thomee et al., 2016). The first row, “DPT only” shows the performance of a network with the frozen DPT depth feature extractor (Ranftl et al., 2021) and a convolutional head. The second row, “ViT only”, demonstrates the performance of only using a small ViT network trained from scratch (Dosovitskiy et al., 2021). The third row, “DPT & ViT in series”, uses the frozen DPT feature extractor followed by a small ViT network. The fourth row, “DPT & ViT in parallel”, represents our chosen architecture depicted in Fig. 3, which utilizes frozen features from the depth network and task-specific features from a separate ViT image encoder in parallel.
We observe that “DPT & ViT in series” exhibits similar performance to “DPT only”, suggesting that adding more computation on top of the frozen DPT feature in addition to our small convolutional head does not provide extra benefits despite the increased computation. When comparing “DPT & ViT in series” with “DPT & ViT in parallel”, we observe an accuracy difference of 14.6 percentage points. This indicates that accurate focal length prediction requires extra task-specific knowledge in addition to depth information. Furthermore, “DPT & ViT in parallel” outperforms “ViT only”, which highlights the importance of features from the pretrained depth network for obtaining a high-performing focal length estimator.
@@ Appendix C Implementation, Training and Evaluation Details
C.1 Datasets
@@
C.3 Baselines
Below we provide further detail on the setup of the baselines.
@@ C.3 Baselines
C.4 Evaluation setup
In evaluating our approach and baselines, we found the range of valid depth values, the depth map resolution used for computing metrics, the resizing approach used for matching the resolution of the ground truth depth maps, and the choice of intrinsics to affect results, sometimes strongly. This is why we made an effort to setup and evaluate each baseline in the same, fair evaluation set up, which we detail below.
@@ Appendix D Applications
Depth-conditioned image synthesis. In this application we stylize an image through a text prompt via ControlNet (Zhang et al., 2023b). To retain the structure of the input image, we predict a depth map from the input image and use it for conditioning the image synthesis through a pre-trained depth-to-image ControlNet SD 1.5 model. Figure 11 shows the input image, prompt, and predicted depth maps and synthesis results for Depth Pro, Deoth Anything v2, Marigold, and Metric3D v2. We find that only Depth Pro accurately predicts the cables and sky region, resulting in a stylized image that retains the structure of the input image. Baselines either miss cables, causing the cable car to float mid-air (Depth Anything v2), or add a gradient to the sky (Marigold).
@@
Synthetic depth of field. Synthetic depth of field can be used to highlight the primary subject in a photo by deliberately blurring the surrounding areas. BokehMe (Peng et al., 2022a) introduces a hybrid rendering framework that marries a neural renderer with a classical physically motivated renderer. This framework takes a single image along with a depth map as input. In this context, it is essential for the depth map to delineate objects well, such that the photo’s subject is kept correctly in focus while other content is correctly blurred out. Furthermore, the depth map should correctly trace out the details of the subject, to deep these (and only these) details correctly in focus. Figure 12 shows the advantage afforded by Depth Pro in this application. (We keep the most salient object in focus by setting the refocused disparity (disp_focus) hyperparameter of BokehMe as the disparity of the object.)

### v2

@@ B.2 Network backbone
B.3 High-resolution alternatives
We further evaluate alternative high-resolution 15361536 network structures and different pretrained weights (Tab. 9). To do this, we test generalization accuracy by training on a train split of some datasets and testing on a val or test split of other datasets, following the Stage 1 protocol for all models in accordance with Tab. 16 and Tab. 17. All ViT models use a patch size of 1616. For weights pretrained with a patch size of 1414 we apply bicubic interpolation to the weights of the convolutional patch embedding layer and scale these weights inversely to the number of pixels (i.e., the weights are reduced by a factor of 1.3). All ViT models use resolution 15361536, for this we apply bicubic interpolation to positional embeddings prior to training. The Depth Pro approach in all cases uses ViT with resolution 384384 and patch size 1616 for both the patch encoder and the image encoder. SWINv2 and convolutional models are pretrained on ImageNet (Deng et al., 2009). Other models use different pretraining approaches described in their papers: CLIP (Radford et al., 2021), MAE (He et al., 2022), BeiTv2 (Peng et al., 2022b), and DINOv2 (Oquab et al., 2024). For the Segment Anything model we use publicly available pretrained weights, which were initialized using MAE pretraining (He et al., 2022) and subsequently trained for segmentation as described in their paper (Kirillov et al., 2023).
@@ B.3 High-resolution alternatives
Importantly, our proposed architecture performs significantly better than straight-forward scaling up the ViT architecture with DINOv2 pretraining. On DIS5K for instance, our architecture improves the boundary recall by relative over DINOv2.
B.4 Depth representation
To assess the effect of the predicted depth representation, we train a ViT encoder and DPT decoder on Hypersim to predict inverse depth, log-depth, or depth. All configurations in this experiment are supervised with a only mean absolute error. Tab. 10 lists the error computed over several depth ranges. We find that supervising the network on depth directly leads to worse results than log-depth or inverse depth. Overall, predicting inverse depth works best, with the largest difference in the regions close to the camera. This makes inverse depth the representation of choice for downstream tasks like novel view synthesis, which benefit particularly from higher accuracy close to the camera.
Hypersim | ||||||
Training objective | 0-1m | 1-2m | 2-4m | 4-8m | 8-16m | 16m |
Inverse-depth | 0.730 | 0.833 | 0.896 | 0.921 | 0.922 | 0.922 |
Log-depth | 0.700 | 0.807 | 0.892 | 0.919 | 0.920 | 0.920 |
Depth | 0.657 | 0.716 | 0.819 | 0.853 | 0.850 | 0.850 |
B.5 Training objectives
@@
HRWSI | Hypersim | Apolloscape | |||||||
Cond. | AbsRel | Log | AbsRel | F1 | Log | AbsRel | |||
1A | 0.166 | 82.1 | 0.083 | 0.259 | 75.4 | 0.221 | 0.156 | 0.339 | 45.6 |
1D | 0.138 | 85.1 | 0.077 | 0.246 | 78.4 | 0.391 | 0.128 | 0.424 | 60.6 |
1B | 0.156 | 83.3 | 0.078 | 0.249 | 77.3 | 0.388 | 0.152 | 0.300 | 47.3 |
1C | 0.150 | 83.7 | 0.074 | 0.235 | 79.9 | 0.442 | 0.084 | 0.235 | 75.6 |
Stage 1 training objectives. We first evaluate loss combinations for the first stage and report results in Tab. 11. Condition 1A only applies a mean absolute error loss to all datasets. For non-metric datasets, we use the scale-and-shift-invariant version. Condition 1B adds gradient losses to all synthetic datasets. We again use the scale-and-shift-invariant version for non-metric datasets. Following our observations from Sec. 3, we propose to apply an appropriate mean absolute error loss as in other conditions depending on a dataset being metric, but apply a scale-and-shift-invariant gradient loss irrespective of a dataset being metric or not (C). We find that loss combinations minimizing gradients (1B, 1C, 1D) consistently outperform just applying an absolute error (1A). Besides improving relative and metric depth estimates, they strongly improve boundary metrics, here up to a factor of 2. Interestingly, minimizing gradients on all datasets outperforms minimizing gradients on just synthetic data. This suggests that the added diversity from real-world datasets more than balances out their potentially noisy ground truth, even for minimizing gradients, which emphasize the noise. The best performance however, is achieved by applying a scale-and-shift-invariant gradient loss on the synthetic datasets (1C). We found that this setting improves convergence and overall performs best.
HRWSI | Hypersim | Apolloscape | |||||||||||
@@
AbsRel | Log | AbsRel | F1 | Log | AbsRel | |||
2A | ✓ | ✓ | ✓ | ✓ | 0.149 | 83.6 | 0.072 | 0.235 | 81.3 | 0.465 | 0.092 | 0.303 | 72.9 |
2B | ✓ | ✓ | ✓ | 0.148 | 83.7 | 0.072 | 0.230 | 81.0 | 0.463 | 0.092 | 0.299 | 73.1 | |
2C | ✓ | ✓ | 0.150 | 83.7 | 0.072 | 0.235 | 80.8 | 0.468 | 0.091 | 0.300 | 73.2 | ||
2D | ✓ | 0.150 | 83.4 | 0.074 | 0.239 | 79.8 | 0.461 | 0.096 | 0.349 | 72.8 | |||
2E | 0.159 | 82.7 | 0.074 | 0.242 | 80.6 | 0.459 | 0.096 | 0.346 | 73.3 |
Stage 2 training objectives. The second stage of our training curriculum focuses on sharpening depth boundaries while retaining high metric depth accuracy. To that end, we only employ synthetic datasets due to their high quality ground truth. The obvious strategy for sharpening predictions is the application of gradient losses. We evaluate our combination of multi-scale derivative-based losses in an ablation study. Condition 2A uses all of the losses, namely , ,, , and . See Tab. 12. 2B removes the second-order loss . 2C further removes the squared first order losses . 2D removes all derivative-based losses. 2E applies the to all datasets. Removing improves results on Apolloscape. Our combination of 0th- to 2nd-order derivative losses (2A) performs best across metrics and datasets in aggregate (e.g., in terms of the average rank across metrics).
B.6 Full curricula
HRWSI | Hypersim | Apolloscape | |||||||
Cond. | AbsRel | Log | AbsRel | F1 | Log | AbsRel | |||
3A (Ours) | 0.149 | 83.6 | 0.072 | 0.235 | 81.3 | 0.465 | 0.092 | 0.303 | 72.9 |
3B | 0.148 | 83.9 | 0.073 | 0.245 | 81.3 | 0.478 | 0.095 | 0.292 | 72.1 |
3C | 0.153 | 83.6 | 0.166 | 0.386 | 37.1 | 0.095 | 0.586 | 0.712 | 0.5 |
We assess the efficacy of our complete training curriculum in comparison to alternatives. Condition 3A represents our two-stage curriculum. Condition 3B trains in a single stage and applies all the second-stage gradient losses throughout the whole training. Condition 3C reverses our two stages and represents the established strategy of pretraining on synthetic data first and fine-tuning with real-world data. We find that training with a single stage is a reasonable default strategy and works much better than first training on synthetic data and then fine-tuning on real data. Our proposed strategy however, yields further improvements on metric depth with slightly worse boundary accuracy.
B.7 Focal length estimation
@@
Encoder for depth only | 60.0 |
Encoder for focal length only | 74.4 |
Encoder for depth and refinement network | 63.6 |
Parallel encoders for depth and focal length | 78.2 |
Controlled evaluation of network structures. We evaluate a number of choices for the focal length estimation head and report results in Tab. 14. The models are evaluated on 500 images randomly sampled from Flickr (Thomee et al., 2016). As the first condition, we extract features from the frozen image encoder trained for depth estimation and add a small convolutional head. As the second condition, we train a separate ViT-based image encoder (Dosovitskiy et al., 2021). As the third condition, we train a ViT-based encoder on extracted features from the frozen image encoder for depth estimation. The final condition represents our chosen architecture depicted in Fig. 3, which utilizes frozen features from the depth network and task-specific features from a separate ViT image encoder in parallel.
We observe that refining depth features performs on par with just using the frozen depth features, suggesting that adding more computation on top of the frozen DPT features in addition to our small convolutional head does not provide extra benefits despite the increased computation. Training a separate image encoder from scratch improves performance by 14.6 percentage points, which indicates that accurate focal length prediction requires extra task-specific knowledge in addition to depth information. Furthermore, the two encoders in parallel outperform using just a single image encoder for focal length prediction, which highlights the importance of features from the pretrained depth network for obtaining a high-performing focal length estimator.
@@ Appendix C Implementation, Training and Evaluation Details
C.1 Merge operation
We merge overlapping feature patches to feature maps by generating a Voronoi partition of the desired feature map. To generate the partition, we use the patch centers as seeds and obtain a Voronoi cell per feature patch. The area of the patch covered by the Voronoi cell is copied to the feature map, the remaining area discarded. By overlapping patches we ensure that the receptive field of the patch encoder partially covers neighboring patches. Fig. 11 illustrates the approach for merging patches into the Feature 4 feature map.
C.2 Focal length head
The focal length head consists of a three-layer convolutional network with kernel sizes 3,3,6, and strides 2,2,1. Each layer halves the channel dimension and is followed by a rectified linear unit. The first layer starts from 128 channels and the last reduces channels to a single focal length value per image.
C.3 Datasets
@@ C.4 Training hyperparameters
C.5 Baselines
Below we provide further details on the setup of the baselines.
@@ C.5 Baselines
C.6 Evaluation setup
In evaluating our approach and baselines, we found the range of valid depth values, the depth map resolution used for computing metrics, the resizing approach used for matching the resolution of the ground truth depth maps, and the choice of intrinsics to affect results, sometimes strongly. This is why we made an effort to set up and evaluate each baseline in the same fair evaluation setup, which we detail below.
@@ Appendix D Applications
Depth-conditioned image synthesis. In this application we stylize an image through a text prompt via ControlNet (Zhang et al., 2023b). To retain the structure of the input image, we predict a depth map from the input image and use it for conditioning the image synthesis through a pretrained depth-to-image ControlNet SD 1.5 model. Figure 12 shows the input image, prompt, and predicted depth maps and synthesis results for Depth Pro, Deoth Anything v2, Marigold, and Metric3D v2. We find that only Depth Pro accurately predicts the cables and sky region, resulting in a stylized image that retains the structure of the input image. Baselines either miss cables, causing the cable car to float in mid-air (Depth Anything v2), or add a gradient to the sky (Marigold).
@@
Synthetic depth of field. Synthetic depth of field can be used to highlight the primary subject in a photo by deliberately blurring the surrounding areas. BokehMe (Peng et al., 2022a) introduces a hybrid rendering framework that marries a neural renderer with a classical physically motivated renderer. This framework takes a single image along with a depth map as input. In this context, it is essential for the depth map to delineate objects well, such that the photo’s subject is kept correctly in focus while other content is correctly blurred out. Furthermore, the depth map should correctly trace out the details of the subject, to keep these (and only these) details correctly in focus. Figure 13 shows the advantage afforded by Depth Pro in this application. (We keep the most salient object in focus by setting the refocused disparity (disp_focus) hyperparameter of BokehMe as the disparity of the object.)